### v1

# Debias your Large Multi-Modal Model at Test-Time with Non-Contrastive Visual Attribute Steering

Abstract
Large Multi-Modal Models (LMMs) have demonstrated impressive capabilities as general-purpose chatbots that can engage in conversations about a provided input, such as an image. However, their responses are influenced by societal biases present in their training datasets, leading to undesirable differences in how the model responds when presented with images depicting people of different demographics. In this work, we propose a novel debiasing framework for LMMs that directly removes biased representations during text generation to decrease outputs related to protected attributes, or even representing them internally. Our proposed method is training-free; given a single image and a list of target attributes, we can ablate the corresponding representations with just one step of gradient descent on the image itself. Our experiments show that not only can we can minimize the propensity of LMMs to generate text related to protected attributes, but we can improve sentiment and even simply use synthetic data to inform the ablation while retaining language modeling capabilities on real data such as COCO or FACET. Furthermore, we find the resulting generations from a debiased LMM exhibit similar accuracy as a baseline biased model, showing that debiasing effects can be achieved without sacrificing model performance.
1 Introduction
Deep neural networks are well known to exhibit a myriad of societal biases learned from their training datasets [8, 45]. While this effect is true to some extent for most models, it is especially prevalent in models trained on web-scale data. Numerous prior works have observed and identified such biases present in modern Large Language Models (LLMs) [5, 9], while other recent work shows that such biases are even more prevalent in Large Multimodal Models (LMMs)  [^1] LMMs refer specifcally to generative multimodal models. We refer to the general class of vision-language models as VLMs [7] such as LLaVA-1.5 [24], or Llama-3.2 [12] that integrate a vision backbone with a pretrained LLM. Given that LLMs are often pretrained on relatively uncurated web-scale data [32], any further integration with additional models will inherit the particular biases of the LLM. Without additional safety tuning, these pre-existing biases may be amplified further when an LLM is augmented with pretrained visual capabilities, which also come with a distinct set of implicit societal biases present in the visual pretraining data. Evaluating and mitigating potentially harmful behaviors induced by these societal biases is becoming increasingly important in order to safely deploy multimodal generative AI systems that utilize LMMs. In this work we focus on reducing the propensity of an LMM to attend to protected attributes unnecessarily. While it is often benign and even useful to make use of protected attributes in generation, LMMs have been found to suffer from group-dependent sentiment [27], well as conditional inaccuracies given certain protected attributes [18, 44].
Figure 1: An overview of our model steering approach for LMMs. By solving for an appropriate steering vector we can augment the generation patterns of the LMM by ablating the internal representation corresponding to a given attribute at test time. The steering vector is computed with a single step of gradient optimization to maximize the likelihood of protected tokens. The attribute that we ablate is then computed by taking a difference between original activations and optimized activations.
Recently, a variety of methods have been proposed for debiasing LLMs and Discriminative Vision Language Models individually [22, 34]. However, relatively little prior work has focused specifically on debiasing LMMs. One naive approach could be to simply define a list of words or symbols that are blocked at inference time. While a block-list could have merit in some applications, a more nuanced approach is required for attributes with many or ambiguous meanings, such as perceived race (e.g., black, white) [^2] Wherever we use words such as race or gender throughout this work, we mean perceived race and perceived gender. The labels we used for such social attributes are inherited from prior work and are not a reflection of our own value judgments. See Section 6 for additional discussion.. Many other existing debiasing approaches for LLMs and VLMs focus on fine-tuning models with additional curated data to reduce bias. However, training is a blunt tool that often results in undesirable outcomes such as a degradation in task-specific performance. Fine-tuning is also labor and computationally intensive, requiring the collection of an additional (likely large-scale) dataset that can appropriately debias the model. Despite prior efforts [17], there remains no cannonical recipe for constructing a dataset large enough to meet data requirements for training LMMs, while also fine-grained enough to ablate specific target attributes like perceived race. In contrast, our work introduces a training-free approach to debiasing LMMs that can be applied to any attribute at inference time without degrading model performance.
Inspired by mechanistic interpretability and prior work in steering language models, we propose a novel method for identifying and ablating biased representations present in LMMs. Recently, prior work has shown that it is possible to ablate representations with respect to target attributes by using a small number of contrastive prompts [26] in order to identify directions in representation space that correspond to an increased propensity to generate related outputs. Our method differs from prior LLM steering work in two fundamental ways. First, our method targets multimodal models rather than just language models. Second, our approach requires no additional data other than an input image and a target attribute to remove. We find that we can leverage the additional information available to a LMM, the input image, in order to solve for the steering direction on the fly. Specifically, during inference we perform a single step of gradient descent on the visual representation in the direction of the target attributes. This one-step optimization is similar to the fast sign gradient method (FGSM) [15] applied on LMM representations. We then can perform model steering as normal, by extracting downstream activations that follow from both the perturbed and the unchanged visual representations. The normalized difference between them is used as the steering direction.
Contributions
Our contributions are three-fold:
  • We propose an efficient training-free algorithm to find effective steering directions for LMMs that performs a single optimization step on the input visual representation.
  • We show that our proposed approach can effectively debias two popular LMM model families LLaVA and Llama3. We reduce their propensity to generate text related to protected attributes on the SocialCounterfactuals, COCO, and FACET datasets. We find that our results hold across perceived race, physical appearance, age, and gender attributes.
  • We analyze the steered and unsteered generations across all datasets and find that steering does not degrade accuracy or language modeling capabilities, finding instead, steering produces equitable group-level sentiment.
2 Related Work
Model Steering via Interventions. Due to the expense and difficulty of fine-tuning LLMs for arbitrary domains, its natural to seek an efficient and fine-grained approach for controlling LLM outputs. Prior work has attempted to steer generations by intervening on inputs, outputs, or intermedate representations. Modifying prompts to induce or ablate a feature [41, 40] is one such approach, but is commonly an iterative and inefficient method. Intervening on outputs has been explored via token-level block lists [14], domain classifiers [21] to trigger regeneration, or decoding strategies that decrease probabilities of generating certain tokens [31]. Intervening on intermediate representations is most relevant to our approach. This often entails the discovery of interpretable features that not only explain model behavior [47, 19], but can also be used to intervene and steer the model towards output generations with certain characteristics or content.
Templeton et al. [36] showed that this can be achieved by applying a sparse autoencoder to decompose the activations of an LLM into separable features. There, the authors demonstrated the existence of monosemantic features that can trigger relevant downstream behavior or content when manually introduced during inference. Turner et al. [38] showed LLMs can be steered with a simple contrastive prompt using vector addition, and Arditi et al. [3] demonstrated that the refusal behavior in LLMs can be suppressed through a single vector which can be learned via ablation on representative data and applying a difference-in-means [4] approach. Various methods of steering have also been applied to toxiciy [25] and other behaviors such as hallucination [28].
Social bias mitigation.
While several approaches have been proposed for mitigating social biases in VLMs [39, 6, 43, 33, 11, 35, 18], prior research on addressing such biases in LMMs is lacking. Sathe et al. [30] and Fraser and Kiritchenko [13] utilized synthetically generated images to analyze the presence of bias in LMMs, but do not address bias mitigation strategies. Howard et al. [17] also leveraged synthetic images from the SocialCounterfactuals dataset [18] to measure bias in LMMs but at a much larger scale, finding that LMMs possess more bias than the corresponding LLM from which they were trained. They also investigated the usefulness of prompting strategies to reduce bias at inference time, but found that it produced inconsistent debiasing effects across different models and generation settings. While representaiton-based steering for reducing societal biases has been demonstrated in LLMs such as Claude 3 [36], our work is the first to demonstrate successful inference-time representation steering for reducing bias in LMMs.
Figure 2: A detailed description depicting our single step gradient-based steering direction computation. (Step 1) Compute the intermediate activations of the LLM when asking the model a direct question about the target attribute to remove. (Step 2) Run a single step of gradient optimization with a logit maximizing loss function in order to compute a modified version of the image tokens $\mathbf{u}$ which increases the probability the model will reference the protected attribute. (Step 3) Recompute the activations of the LLM with the new modified image tokens $\mathbf{u^{\prime}}$ and take a difference between the two intermediate LLM activations. (Step 4) Run the intended query for the LMM, but ablate the attribute $a$ from all intermediate layers of the LLM.
3 Methodology
Our approach to debiasing LMMs involves identifying and ablating the bias attribute in the model’s internal representations. We achieve this by contrasting the model’s original activations from alterated activations: either from a single step gradient optimized version or from a collection of bias invoking questions. An overview of the single step gradient-based method is shown in figure 2.
3.1 Steering Direction Estimation
Dataset-based Estimation
Let $\mathcal{M}$ denote an arbitrary LMM, and $\mathbf{h}^{l}\in\mathbb{R}^{d}$ represent the activations at layer $l$, where $d$ is the dimensionality of the hidden state. We use $\mathbf{a}\in\mathbb{R}^{d}$ to denote the bias attribute, which is a vector that captures the direction of the bias in the model’s internal representations, and define $\mathbf{r}^{l}$ as the residual at layer $l$.
To estimate the bias attribute, we collect a dataset of standard prompt-image pairs $\mathcal{D}_{\text{standard}}=\{(\mathbf{x}_{i})=(\mathbf{p}_{i},\mathbf{i}_{i})\}_{i=1}^{N_{\text{standard}}}$ and a dataset of prompt-image pairs which elicit biased responses $\mathcal{D}_{\text{bias}}=\{(\mathbf{x}_{i})=(\mathbf{p}_{i},\mathbf{i}_{i})\}_{i=1}^{N_{\text{bias}}}$. Here, $\mathbf{p}_{i}$ represents the text prompt and $\mathbf{i}_{i}$ represents the corresponding image. We compute the activations of the model on both datasets and calculate the difference in means:
$\mathbf{a}=\frac{1}{|\mathcal{D}_{\text{bias}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{bias}}}\mathbf{h}^{l}(\mathbf{x})-\frac{1}{|\mathcal{D}_{\text{standard}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{standard}}}\mathbf{h}^{l}(\mathbf{x})$
Single Step Gradient-based Estimation
By having access to the models weights and a list of words related to a target attribute, we can perform a single step of gradient optimization to create a contrasting set of inputs to the language model. We define $T$ be a list of tokens related to a target attribute.We overload $h^{l}(x,u)$ to represent the output of layer $l$ given the input text $x$ and encoded image tokens $u$.
First, we run a single step of gradient computation on the model to encourage it to mention protected words.
$\mathcal{L}_{\text{bias}}=\frac{1}{|T|}\sum_{t}^{T}||(logit_{t}-(M_{t}))||_{2}^{2}$
where $M_{t}=logit_{t}+max(logits)$. The choice of loss function $L_{bias}$ is discussed and analyzed in section 5.2. Next we compute the gradient w.r.t. the encoded image tokens and simply use the sign of the gradients to modify the image tokens.
$\mathbf{u}^{\prime}=\mathbf{u}+\epsilon\cdot\text{sign}(\nabla_{\mathbf{u}}\mathcal{L}_{\text{bias}})$
where $\epsilon$ is a small constant, and we find computing it based on the average token magnitude to work consistently $\epsilon=\frac{\sigma(|\mathbf{u}|)}{2}$ with $\sigma(\mathbf{u})$ denoting the standard deviation of the image tokens. Then we can simply compute the attribute vector with the modified image representation to compute the linear direction corresponding to the attribute: $\mathbf{a}=\mathbf{h}^{l}(\mathbf{x,u^{\prime}})-\mathbf{h}^{l}(\mathbf{x,u})$
Test-Time Steering
We normalize the bias attribute to have unit length: $\mathbf{a}\leftarrow\mathbf{a}/\|\mathbf{a}\|_{2}$. To ablate the bias attribute, we project the residual at each layer onto the bias attribute and subtract the projection from the residual to get a new residual $\mathbf{r}^{l^{\prime}}=\mathbf{r}^{l}-\mathbf{a}\mathbf{a}^{\top}\mathbf{r}^{l}$. We apply this ablation process to every residual in the LMM, effectively removing the bias attribute direction from the model’s internal representations.
3.2 Evaluating Steered Models
Following the approach of Howard et al. [17], we evaluate the effectiveness of our steering methods for debiasing through large-scale generation of LMM responses to open-ended prompts. Specifically, given an input image, we generate responses to prompts such as Describe this image in as much detail as possible and What are 5 keywords that describe the characteristics of this person? We use a total of five such prompts inherited from prior work [17] and produce three different LMM responses to each image and prompt by varying the random seed. Our main goal with evaluating our models is to directly measure the effects on bias in an open ended setting.
3.2.1 Datasets
In order to evaluate how open-ended LMM generations vary in response to images depicting people of different social attributes, we utilize images from three datasets which possess social attribute annotations: DA-COCO [10, 44], SocialCounterfactuals [18], and FACET [16]. The SocialCounterfactuals dataset contains 171k synthetic images depicting people of various races, genders, and physical characteristics such as age and body type. Images are grouped into counterfactual sets, in which each image is highly similar in its depiction of a common subject (e.g., a doctor) while varying only according to intersectional social attributes (e.g., race & gender). We use separate subsets of the SocialCounterfactuals dataset for constructing ablation directions and evaluating models for bias related to race, gender, age, and body type. The Demographic Annotations on COCO (DA-COCO) dataset contains 28k images of people with annotations for skin tone and gender, which we use for evaluating bias related to perceived race and gender. Finally, the FACET dataset contains 32,000 diverse, high-resolution images annotated by experts, covering 13 person attributes and 52 person classes. For our evaluation, we utilize the perceived age, race, and gender attribute labels.
3.2.2 Synthetic Image Generation
An essential component of our method is that we can compute steering directions for LMMs from fully synthetic data, that can be successfully applied to real data. For each target attribute e.g. perceived race, age, physical appearance, and gender, we use synthetically generated images from two sources to compute steering directions: the SocialCounterfactuals dataset, and images generated with the Flux-1.0-dev image generator [1]. In the case of SocialCounterfactuals, the images are generated with Stable Diffusion [29], and are constrained to be counterfactual with respect to the target attribute in question e.g. perceived race. For Flux, images align with the target protected attributes but are not constrained to be counterfactuals.
3.2.3 Protected Attribute Mentions
We measure the frequency at which LMMs mention an individual’s protected attributes during open-ended generation, which can be undesirable in circumstances in which commenting on a person’s race, age, gender, or body type would be considered offensive or inappropriate. While our evaluation method can be applied to any social attribute that relates to a set of target tokens, our evaluations focus on four protected attributes of perceived race, age, gender, body type. Due to the variety of ways in which such attributes can be referenced in natural language, we employ two alternative approaches for measuring their frequency: bigram counts and LLM-as-a-judge.
Bigram frequencies.
We define a list of target words related to the attribute in question and detect all bigrams in model generations beginning with these words. Since many attribute-related terms are polysemous, we hand-annotate the most frequent bigrams to filter out unrelated terms. This enables us to adjust for over- or under-counting by including only those bigrams that have been verified as attribute-related or excluding those that have been annotated as unrelated. Despite being transparent and interpretable, this approach is limited in the sense that it may not capture all possible ways in which protected attributes could be referenced.
SocialCounterfactuals FACET DA-COCO
Model Method Evaluator Age Gender Race Body Type Age Gender Race Gender Race
LLaVA-1.5 Unsteered Bigram 595 358 790 111 232 439 635 617 374
LLaVA-1.5 Steered Dataset Bigram 197 483 549 31 70 282 543 537 274
LLaVA-1.5 Steered Gradient Bigram 583 343 784 109 235 421 650 598 374
Llama-3.2 Unsteered Bigram 507 233 1480 101 138 329 2537 607 1824
Llama-3.2 Steered Dataset Bigram 526 130 629 24 76 120 635 329 456
Llama-3.2 Steered Gradient Bigram 474 190 1240 65 118 337 2380 670 1805
LLaVA-1.5 Unsteered GPT-4 771 1103 235 329 242 1095 21 2689 82
LLaVA-1.5 Steered Dataset GPT-4 365 842 46 158 102 949 1 2269 17
LLaVA-1.5 Steered Gradient GPT-4 783 945 257 301 257 1117 23 2672 68
Llama-3.2 Unsteered GPT-4 869 1641 437 259 246 2169 136 3396 202
Llama-3.2 Steered Dataset GPT-4 893 1564 194 121 125 2800 24 3589 69
Llama-3.2 Steered Gradient GPT-4 792 1312 329 198 162 1872 77 3801 163
Table 1: Frequency of protected attribute mentions per 1k generations for our steered models and other baselines when presented with images from different datasets and open-ended prompts. We find that steering with attributes computed from a dataset nearly always reduces the mention of protected attributes w
LLM-as-a-judge.
We additionally use GPT-4o [2] as a judge to annotate the amount of attribute-related text in each generation. Using a two-shot prompt with OpenAI’s Structured Output API, GPT-4o returns both the count of protected attributes and their corresponding spans. This method has proven to be highly reliable, with minimal discrepancies between the reported counts and the identified spans. Our manual inspection of GPT-4o’s highlighted spans confirmed that it captures a broad, but justified, set of terms that refer to the target attributes.
3.2.4 Sentiment analysis
Another potential form of bias in LMMs is a disparity across groups in the sentiment of generated text sequences. To investigate whether our steering approach can reduce such disparities, we use VADER [20], a popular lexicon and rule-based sentiment analysis tool [37, 42]. Specifically, we report the compound score produced by VADER across different groups of perceived race, gender, age, and body type attributes in our evaluated datasets. A compound score $>=0.05$ indicate positive sentiment while scores $<=-0.05$ indicate negative sentiment.
3.2.5 Occupation mentions
The SocialCounterfactuals and FACET datasets both contain annotations for the occupation of the individual depicted in each image. Since the person’s occupation is relevant and likely to be mentioned by LMMs in response to our open-ended prompts, we measure the presence of this additional attribute in generated text to evaluate whether our steering has other unintended effects on the quality of generations. Specifically, we count exact matches for the annotated occupation name in each generation produced for images from FACET and SocialCounterfactuals.
4 Experimental Results
4.1 Experimental details
Generation
We conduct experiments by generating responses from two LMMs with and without our steering methods: LLaVA 1.5 [23] and Llama-3.2-Vision-Instruct [12]. We generated a total of 1,820,904 responses for the five prompts, three datasets, two LMMs, and three methods evaluated in our experiments. Our generation experiments were conducted on 512 Intel® Gaudi 2 AI accelerators in the Intel® Tiber AI Cloud.
Selecting a Steering Direction
For each LMM, we compute steering directions by computing the normalized difference between contrasting activations from biased and benign text generations. Biased text is generated from a specific prompt applied to 256 image samples, while benign text is sourced from the LLaVA-Instruct-80K dataset [24] by excluding instances with protected attributes. We evaluate 32 candidate ablation directions based on a held-out set of 10 image-prompt pairs, selecting the most effective direction for further experiments.
4.2 Protected attribute evaluation
Table 1 shows our main results for evaluating the effectiveness of steering LMM generations from mentioning protected attributes. Across all three datasets, we see that on average and across nearly all conditions, steering significantly reduces the LMM’s propensity to generate text related to the target attributes. This result holds for both the LLM-as-judge as well as the bigram count evaluation methods. One clear trend is that as the model capability increases, so does the effectiveness of our Steered Gradient steering approach. While Steered Dataset is nearly always more effective for the 7B LLaVA model, the optimization-based method is more viable for Llama-3.2-11B. What’s more, Steered Gradient is an improvement over Steered Dataset in the few cases where Steered Dataset shows no improvement over the unsteered baseline. As mention frequency of protected attributes is the most straightforward metric on our approach for debiasing, we believe this shows a significant improvement over the baseline, with an overall improvement of 29% for our Steered Dataset method.
4.3 Group-specific sentiment
Model Method Variance Range
LLaVA-1.5 Unsteered 57.6 20
LLaVA-1.5 Steered Dataset 41.6 14
LLaVA-1.5 Steered Gradient 16.4 11
Llama-3.2 Unsteered 197.9 41
Llama-3.2 Steered Dataset 169.2 29
Llama-3.2 Steered Gradient 309.9 49
Table 2: Variance and range of negative sentiment scores per 1k generations, measured across perceived race groups using the SocialCounterfactuals dataset.
To investigate the impact of our steering methods on differences in the sentiment of text produced for different protected attribute groups, we calculated the frequency of negative text detections per 1k generations for each perceived race group annotated in the SocialCounterfactuals dataset. Table 2 provides the variance and range of the resulting values, calculated across race groups. Higher values of variance and range indicate a greater degree of variability in the frequency of negative text generation across different groups. We observe that both of our steering approaches achieve reductions in this measure of bias with LLaVA-1.5, with our Steered Gradient method providing a particularly large reduction in negative sentiment variability. Additionally, our Steered Dataset approach also reduced negative sentiment variability for Llama-3.2. These results point to additional debiasing effects produced by our steering method beyond simply reductions in the frequency of protected attribute mentions.
4.4 Occupation mentions
Model Steering Method SocialCounterfactuals FACET
Age Gender Race Body Type Age Gender Race
LLaVA-1.5 Steered Dataset -19.43% -17.38% -30.12% -13.18% -2.77% -4.50% -18.34%
LLaVA-1.5 Steered Gradient 0.46% -6.23% -0.08% 1.85% -0.69% 0.00% -3.81%
Llama-3.2 Steered Dataset -5.48% -15.02% -7.84% -2.83% 42.37% 38.55% 4.20%
Llama-3.2 Steered Gradient -3.96% -0.40% 2.90% 8.53% -9.86% 7.25% -7.63%
Table 3: Comparison of Steering on open ended accuracy mentions with LLaVA and Llama LMMs relative to their unsteered baselines for SocialCounterfactuals and FACET.
In table 3 we measure how often the steered models correctly mention the occupation of the person in the image in our open ended prompt generation experiments. While we find the occupation mentions often go down, there is a high variance across the experimental settings, with the interesting result for Llama-3.2 on FACET significantly increasing occupation mentions. We also find the Steered Gradient method to stay more similar in its overall generations to the baseline, hence its lower changes with respect to the unsteered models.
4.5 Accuracy of steered models
We employed the LLM-as-a-judge approach [46] to investigate whether steering affects the accuracy of generated responses. We used GPT-4o to evaluate whether LLaVA’s text responses, with and without steering, match the corresponding image. GPT-4o was given the image and prompt: “Does the description match the image? Answer with Yes or No.” Manual analysis showed that GPT-4o responds “No” when the generation contains extra details not present in the image. The results show no significant difference in accuracy between baseline and steered LLaVA models, indicating that steering does not degrade performance.
5 Analysis
5.1 Impact of steering on token probabilities
Figure 3: The global changes in probabilities of predicting given tokens on the age subset of FACET. The table reflects generated outputs from 100 samples, sorted by most changed, showing how overall steering only effects the raw probabilties in a sublte way.
Figure 3 shows the effectiveness of steering techniques in reducing bias in LMM token predictions. After intervening to ablate biased directions in the model’s internal representations, we observe a shift toward more neutral, contextually appropriate tokens, with biased terms related to protected attributes being suppressed. This effect is consistent in both single-image examples and across 100 generations from the FACET age split.
5.2 Loss Function Analysis for One Step Gradient-Based Optimization
We investigated various loss functions to determine their suitability for gradient-based token manipulation. While many intuitive loss functions appeared promising, we found that they often yielded ineffective results. For instance, using the sigmoid activation on the logits led to nearly zero gradients in practice, as the logit values in the LLM are often high-magnitude, saturating the sigmoid function and nullifying gradients.
To better understand the behavior of different loss functions in the context of our proposed feature optimization, we analyze the gradients of three key functions: KL Divergence, repeated Cross-Entropy, and Mean Squared Error. Each loss function is evaluated in terms of its impact on the gradient direction and magnitude with respect to the input logits of the LLM, denoted as $\mathbf{z}\in\mathbb{R}^{V}$ for a vocabulary size $V$.
Let $\mathbf{p}=\text{softmax}(\mathbf{z})$ represent the probability vector obtained from the LLM logits, and let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{N}\}$ be the set of target tokens we aim to manipulate. We analyze the gradients of the following loss functions: KL-Divergence (KL), a column-wise Cross-Entropy (CE), and Mean Squared Error (MSE), to understand their impact on the feature optimization.
Maximizing Target Token Probabilities
: Here we define the target distribution $\mathbf{q}$ as uniform over the target set $\mathcal{T}$, assigning equal probability to each token $i\in\mathcal{T}$:
$q_{i}=\begin{cases}\frac{1}{N}&\text{if }i\in\mathcal{T}\\ 0&\text{otherwise}\end{cases}$
The KL-Divergence between the target distribution $\mathbf{q}$ and the model’s predicted probability distribution $\mathbf{p}$ is:
$\text{KL}(\mathbf{q}||\mathbf{p})=\sum_{i=1}^{V}q_{i}\log\frac{q_{i}}{p_{i}}=\sum_{i\in\mathcal{T}}\frac{1}{N}\log\frac{1/N}{p_{i}}$
The gradient of this loss function with respect to each logit $z_{j}$ is computed as:
$\frac{\partial}{\partial z_{j}}\text{KL}(\mathbf{q}||\mathbf{p})=-\sum_{i\in\mathcal{T}}\frac{1}{N}\frac{1}{p_{i}}\frac{\partial p_{i}}{\partial z_{j}}$
where the gradient of the resulting softmax $\frac{\partial p_{i}}{\partial z_{j}}=p_{i}(\delta_{ij}-p_{j})$ yields:
$\frac{\partial}{\partial z_{j}}\text{KL}(\mathbf{q}||\mathbf{p})=-\sum_{i\in\mathcal{T}}\frac{1}{N}(\delta_{ij}-p_{j})$
This results in a cumulative gradient heavily influenced by $-p_{j}$, yielding a relatively small gradient magnitude when $p_{j}$ values are spread close to uniform over $\mathcal{T}$. Consequently, KL-Divergence provides weak gradients in settings where uniform probability across the target tokens is desired, limiting its effectiveness.
Column Cross-Entropy on Tokens
To shift the probability distribution towards individual tokens in $\mathcal{T}$, we apply a repeated Cross-Entropy (CE) loss for each $t\in\mathcal{T}$:
$\text{CE}(\mathbf{p},\mathcal{T})=-\sum_{t\in\mathcal{T}}\log(p_{t})$
For a single token $t$, the Cross-Entropy gradient with respect to $\mathbf{z}$ is:
$\frac{\partial}{\partial z_{j}}(-\log(p_{t}))=-(\delta_{tj}-p_{j})$
where $\delta_{tj}$ is 1 if $j=t$ and 0 otherwise. Summing over all tokens in $\mathcal{T}$ yields:
$\frac{\partial}{\partial z_{j}}\text{CE}(\mathbf{p},\mathcal{T})=-\sum_{t\in\mathcal{T}}(\delta_{tj}-p_{j})$
This form introduces a “tug-of-war” effect, where each token $t\in\mathcal{T}$ pulls the logits towards itself, creating interference among the tokens in $\mathcal{T}$. The resulting gradient direction is not aligned with maximizing probabilities across all tokens in $\mathcal{T}$ simultaneously, leading to inconsistent results.
Mean Squared Error Loss with Target Logits
Lastly, we minimize the MSE on the target tokens, setting a target logit value $M$ for each token in $\mathcal{T}$:
$\text{MSE}(\mathbf{z},\mathcal{T},M)=\sum_{i\in\mathcal{T}}(z_{i}-M)^{2}$
The gradient of this loss with respect to each logit $z_{j}$ is:
$\frac{\partial}{\partial z_{j}}\text{MSE}(\mathbf{z},\mathcal{T},M)=\begin{cases}2(z_{j}-M)&j\in\mathcal{T}\\ 0&j\notin\mathcal{T}\end{cases}$
This gradient is directly proportional to $z_{j}-M$, ensuring a strong directional push for each $z_{j}$ in $\mathcal{T}$ towards $M$. Additionally, only logits corresponding to tokens in $\mathcal{T}$ are affected, avoiding interference and aligning the gradient direction to effectively manipulate each token’s logit towards the target value $M$. Therefore, we run our Gradient-based Steering with MSE loss.
6 Discussion
Our findings underscore the substantial benefits of a rapid, training-free technique for model steering in large multi-modal models. By using a single gradient step, we effectively guide the model to downplay specific protected attributes, such as perceived age, race, or gender, in its outputs. Given the rapid deployment of LMMs across high-impact sectors, our approach represents a scalable and impactful method for aligning models with ethical and societal expectations without extensive retraining or re-tuning requirements. Despite our best efforts to improve the fairness of generative AI models, we acknowledge that our choice of models, methodologies, and datasets may themselves contain latent biases which limit our ability to address this multi-faceted problem.
The advantages of this model-steering method are twofold: it is not only effective, but also incredibly efficient. Our approach leverages a single gradient step to reduce or remove attribute-based patterns in model outputs. By circumventing traditional training cycles, which can be prohibitively costly and time-consuming, this method achieves a streamlined pathway to targeted model adjustments. The elimination of a need for full model re-training or fine-tuning directly translates to significant savings in both energy and hardware resources. As AI practitioners grapple with the environmental impact of increasingly large models, our method contributes a sustainable alternative. This not only allows for quick model refinement but also aligns with the broader goals of sustainable AI development, where resource-intensive operations can be minimized without sacrificing efficacy.
Limitations
We acknowledge that our work contains statements regarding social attributes such as race, gender, age, and body type used which may be considered harmful or stereotypical. Additionally, the groups considered for each of these social attributes are not exhaustive and cannot fully characterize the unique aspects of an individual’s identity. Our aim is simply to evaluate sources of bias in LMMs and our ability to mitigate them; it is not our intention to impose such labels on any individual. We are constrained by social demographic annotations which are available in the existing resources employed in our experiments; any missing demographic labels are a reflection of these limitations and not of our own value judgments.
The datasets employed in our evaluations as well as our automated evaluation methods which employ LLMs may themselves contain biases which could influence our results. Nevertheless, we feel that our manual validations and the scale of the study we conducted necessitated the use of these resources. We also acknowledge that our study was only conducted with the English language. Additional work is needed to explore the effectiveness of our steering methods for other languages. Despite these limitations, we hope that the findings of our study will raise awareness of the need for methods to debias LMMs and promote the development of more responsible and fair generative AI models.
Future Work
Finally, our single-gradient-step approach offers a promising framework for further development in bias mitigation and model alignment. This initial study suggests a strong foundation for rapid intervention in model behavior, which could be extended to other forms of model control. Given the success observed in attribute suppression, further research may explore other types of steering or alignment tasks that could similarly benefit from a minimally invasive, gradient-based intervention. Next steps should aim to expand bias mitigation techniques to encompass a broader spectrum of attributes and assess the long-term impacts of steering interventions on model performance.